OpenStudy (anonymous):

how do you solve the system 4x + 3y = -1 3x + y = 3 in the elimination method

OpenStudy (anonymous):

\[4x+3y=−1\] \[3x+y=3\] multiply the second equation by 3 you get : \[9x+3y=9\] now you can eliminate the equations

OpenStudy (anonymous):

\[-5x=-10\] you got \[x=2\]

OpenStudy (anonymous):

so you subtract the two problems from each other?

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

do it urself u reterd

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

@Jescy21 but make sure that both of equations are ready for subtraction, okay

OpenStudy (anonymous):

how do I make sure both equations are ready @AntarAzri

OpenStudy (anonymous):

when you can elimiate one of the unknown term by substracting the equations

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

see first response by AntarAzri multiply the second equation by 3 so when you subtract, y got eliminated..right

OpenStudy (anonymous):

as @isag said, i multiplied by 3 to eliminate the 3y